During a period of exercise the oxygen demand increased. it was necessary for the blood to deliver 10cm³ of oxygen per minute per 100g of respiring skeletal muscle.
-oxygenated blood carries 20cm³of oxygen per 100cm³ of blood
-80% of the Cardiac output is directed at skeletal muscles when exercising
-skeletal muscles account for 50%of lean body mass
a.Calculate the volume of blood required to deliver sufficient oxygen to the muscles of a person weighing 108kg during exercise?
b.What is the required cardiac output to supply oxygen at this rate?
c.If the person had a heart rate of 193 bpm during exercise what would their stroke volume have been?

Many thanks
1
Expert's answer
2016-08-26T08:17:03-0400
a. The mass of skeletal muscles: 108 ·0.5 = 54 kg
Oxygen requirement for the muscles: 10 cm³ · (54/0.1) = 5400 cm³ of oxygen.
Oxygenated blood requirement: 5400/ 20 = 270 cm³ of blood.

b. Required cardiac output: 270/0.8=338 cm³ of blood.

c. Heart rate = Cardiac output / stroke volume
Stroke volume = cardiac output / heart rate = 338 / 193 / 60 = 105 cm3.